Find the equation of a line that is parallel to the line whose equation is x+2y=6 and whose y-intercept is -1
Answer by fractalier(2101)

Since parallel lines have equal slopes, we need a line with the same slope as x + 2y = 6...this can be found by solving for y, or even easier, look at m = -A/B...
in x + 2y = 6, that would be m = -1/2.
Now the line we need goes thru (0, -1) and we use either the point-slope method, or even easier, the slope-intercept form...so we have
y = mx + b
y = (-1/2)x - 1